OpenStudy (anonymous):

Look at the table. Polyhedron Number of Faces Number of Edges Number of Vertices Polyhedron 1 4 6 4 Polyhedron 2 6 12 8 Polyhedron 3 8 10 6 Polyhedron 4 12 30 20 Which polyhedron has incorrect entries for its number of faces, edges, or vertices? Answer Polyhedron 1 Polyhedron 2 Polyhedron 3 Polyhedron 4

Directrix (directrix):

Polyhedron 3 --------- 8 = F 10 = E 6 = V --------------- The relationship among the vertices, faces, and edges of a polyhedron is one of Euler's formulas: V + F = E + 2 where V = number of vertices, F = number of faces, and E = number of edges. In polyhedron 3, 6 + 8 = 10 +2 is not true.